The curve has an equation y = e*. Compute the area bounded by the


curve from x = 0 to x = 1.

1
Expert's answer
2022-06-06T23:29:57-0400
"A=\displaystyle\int_{0}^1e^xdx=[e^x]\begin{matrix}
 1 \\
 0
\end{matrix}=e-1\ ({units}^2)"